All of the chlorine in the compound CoCl2 [cobalt (II) chloride] represents what percent of the compound’s molar mass?
1
Expert's answer
2020-12-15T07:52:26-0500

Solution:

"CoCl_2=129.8392( \\tfrac{g}{mole}\t)"

"Cl_2=70.9060( \\tfrac{g}{mole}\t)"


"\\dfrac{70.9060( \\tfrac{g}{mole}\t)}{129.8392( \\tfrac{g}{mole}\t)}\t=0.5461*100\\%=54.6\\%"


Answer:

54.6% of chlorine in the chemical compound [cobalt (II) chloride].